OpenStudy (anonymous):

Can someone help graph this? y=-1/3x+5

OpenStudy (phi):

I vote that katie graph it.

OpenStudy (anonymous):

hahaha noooooooooo never

OpenStudy (anonymous):

This is a linear equation, and it is in the point-slop form: \[y=mx+b\]Where m is the slope and b is the y-intercept.

OpenStudy (phi):

Do you know how to plot a point if you are given its (x,y) coordinates ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

you are told y = (-⅓)x+5 let x=0. in your equation , erase the x and put in 0 remember that -⅓ x is short for -⅓ * x (-⅓ times x) simplify what do you get ?

OpenStudy (anonymous):

y=5?

OpenStudy (phi):

yes, because y = -⅓*0 + 5 y = 0+5 y=5 all of that means x=0, y=5 or (0,5) is on the line. plot that point.

OpenStudy (phi):

next we find another point. we can pick any x we want. (0 is good because the arithmetic is easy). another "easy" point is when x=3 in y = (-⅓)x+5 replace x with 3 and find y y = -⅓ * 3 + 5 can you simplify that ?

OpenStudy (anonymous):

I understand it now, thank you.

OpenStudy (phi):

-⅓ * 3 is the same as -⅓ * 3/1 and we multiply top*top and bottom * bottom \[ \frac{-1}{3}\cdot \frac{3}{1} = \frac{-3}{3} = -1\]

OpenStudy (phi):

you should get when x=3, y= 4 plot (3,4) now draw a straight line through the two points. In theory the line goes on forever in both directions, but you can stop when you get tired.